OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (anonymous):

Distribute the numbers: -2(x+3) = -2(x+1) - 4 -2x -6 = -2x - 2 - 4

OpenStudy (anonymous):

Then you will combine the like terms which is -2 -4 = -6 -2x - 6 = -2x - 6

OpenStudy (anonymous):

Then you will combine like terms again so -2x - 6 = -2x - 6 Add 6 to both sides = 0 and add -2x to both sides which also = 0

OpenStudy (anonymous):

so what would the answer be??

OpenStudy (anonymous):

No real solution :)

OpenStudy (anonymous):

thank :)